OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)? Answer y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (anonymous):

y = 2(x + 2)

OpenStudy (aroub):

slope= 8-0/2+2 =8/4 =2 y-y0=m(x-x0) , m=2 , (2,8) y-8=2(x-2) y-8=2x-4 y=2x+4 y=2(x+2)